Grade 12th passPhysical Chemistry

One gram of metal equivalent weight 40 was burnt in excess of atmospheric oxygen. The metal oxide formed. Was dissolved in 100 ml of water and solution is diluted up to 5 liter. Normality of. Solution is

tulasisai pechetti
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago
Since atomic weight 40 belongs to calcium, hence the given metal is calcium(Ca), the Reaction for the Chemical Process is as follows:

Ca+ 1/2 O₂ =CaO

CaO +H₂O = Ca(OH)₂

The weight of Ca =40

The molar weight of CaO =40 +16 =56

Molar weight of Ca(OH)₂ = 74

equivalent weight of Ca(OH)₂ = 74/2 = 37

40 grams of Ca produces 74 grams of Ca(OH)₂

1 gram of Ca is will produce 74/40 gram Ca(OH)₂ =1.85

Or

1.85/37 = 0.05 gram equivalent of Ca(OH)₂

Therefore 0.05gram is dissolved in 100 ml

Hence normality of the solution = 0.05 x (100/1000) = 0.005 N